OpenStudy (anonymous):

The hypotenuse of a right triangle whose legs are consecutive whole numbers is 29 units. What is the sum of the lengths of the two legs?

OpenStudy (id21):

\[x^2+(x+1)^2=29^2\]

OpenStudy (id21):

find x+x+1

OpenStudy (anonymous):

wait.. what? so the sum of the 2 lengths of the legs are x+x+1?

OpenStudy (anonymous):

2x^2+2x+1 = 841

OpenStudy (anonymous):

thank you

OpenStudy (phi):

The legs are consecutive whole numbers. id21 called the first leg x, and the second leg must be one more than x (that's what consecutive means) the sum of the two legs is x + (x+1), or simply x+x+1 or 2x+1

OpenStudy (id21):

\[x^2+x-420=0\] x=20 (other root is -21 -> you don't need it because it is negative) so the lenght of 1st leg is 20, 2nd leg is 20+1=21, the sum of legs is 20+21=41